# Learning Color Equivariant Representations

### v4

@@ A Hue and saturation group action
In the remainder of this section we refer to Equation 18 as ‘property one’ and Equation 19 as ‘property two’.
A.1 Hue group action
@@
$\varphi_{h}(h_{i},x)=((x_{h}+h_{i})(\text{mod}\;255),x_{s},x_{l}),$ (20)
in Equation 3 of the main text.
@@
$\varphi_{h}(h_{i},\varphi_{h}(h_{j},x))$ $=\varphi_{h}(h_{i},((x_{h}+h_{j})(\text{mod}\;255),x_{s},x_{l}))$ (21)
$=(((x_{h}+h_{j})(\text{mod}\;255)+h_{i})(\text{mod}\;255),x_{s},x_{l}))$ (22)
$=(((x_{h}+(h_{j}+h_{i}))(\text{mod}\;255),x_{s},x_{l}))$ (23)
@@
A.2 Saturation group action
Here we show the proposed saturation group action is a group action. The proposed saturation group action is defined:
@@ A.2 Saturation group action
in Equation 5 of the main text. First we show the proposed saturation group action $\varphi_{s}$ satisfies property one. For any saturation shifts $s_{i}$ and $s_{j}$, we have
@@ B.1 Lifting layer
B.2 Color sorting on CIFAR-10
We present visualizations of the sorted automobile class using Hue-N and CEConv-N architectures. We trained Hue-3, CEConv-3, Hue-4 and CEConv-4 on CIFAR-10 and used the resulting representations to sort the automobile class by color. Figures 9, 10, 11, and 12 show the automobile class sorted by Hue-3 (ours), CEConv-3, Hue-4 (ours), and CEConv-4 respectively. Figures 9, and 10 are qualitatively similar; however, Figures 11 and 12 are not. In particular, notice that there are blue images in the green band (around the center) of Figure 12, but not in Figure 11.
B.3 Hue-shift MNIST feature map visualization
@@
C Datasets
C.1 Hue-shift MNIST
@@
C.2 Hue-shift 3D Shapes
C.3 small NORB
@@ C.4 Tiny-ImageNet
D Implementation and training details
In this section we provide architectural details for our equivariant networks, and training details for each experiment. Our source code is publicly available at https://github.com/CAB-Lab-Princeton/Learning-Color-Equivariant-Representations.
@@ D.2 Hue-shift 3D Shapes
Following (Wiles et al., 2022), we train the ResNet architectures for 100k iterations with a batch size of 128. We optimize over a cross-entropy loss using SGD with a learning rate of $10^{-2}$. Images are down-sampled by a factor of 2 to train the CNN architectures. CNN architectures are trained as described in Section D.1.
D.3 Camelyon17
@@ D.11 ImageNet
E Additional Results
E.1 Luminance shift small NORB
We demonstrate improved generalization to luminance-shifts compared to ResNet18 (He et al., 2016) on the small NORB classification dataset (LeCun et al., 2004). The small NORB dataset consist of 5 categories of objects under different lighting conditions. Variation in images reflects variation in the location and strength of the lighting. Additional details are provided in Appendix C.3.
Networks A/A A/B A/C Params
ResNet18 8.32 (0.86) 37.70 (1.51) 33.88 (5.11) 11.2M
Lum-3* 7.04 (0.65) 18.45 (4.83) 25.14 (1.07) 11.1M
@@ E.1 Luminance shift small NORB
E.2 ImageNet
We include classification results on the ImageNet dataset (Deng et al., 2009). Additional details are provided in Appendix D.
@@ E.2 ImageNet
E.3 Foreground-Background Segmentation

### v6

@@ B Hue and saturation group action
In the remainder of this section we refer to Equation (46) as ‘property one’ and Equation (47) as ‘property two’.
B.1 Hue group action
@@
$\varphi_{h}(h_{i},x)=((x_{h}+h_{i})(\text{mod}\;2\pi),x_{s},x_{l}),$ (48)
in Equation (3) of the main text.
@@
$\varphi_{h}(h_{i},\varphi_{h}(h_{j},x))$ $=\varphi_{h}(h_{i},((x_{h}+h_{j})(\text{mod}\;2\pi),x_{s},x_{l}))$ (49)
$=(((x_{h}+h_{j})(\text{mod}\;2\pi)+h_{i})(\text{mod}\;2\pi),x_{s},x_{l}))$ (50)
$=(((x_{h}+(h_{j}+h_{i}))(\text{mod}\;2\pi),x_{s},x_{l}))$ (51)
@@
B.2 Saturation group action
Here we show the proposed saturation group action is a group action. The proposed saturation group action is defined as
@@ B.2 Saturation group action
in Equation (5) of the main text. First we show the proposed saturation group action $\varphi_{s}$ satisfies property one. For any saturation shifts $s_{i}$ and $s_{j}$, we have
@@ C.1 Lifting layer
C.2 Hue-shift MNIST feature map visualization
@@
C.3 Color sorting on CIFAR-10
We present visualizations of the sorted automobile class using Hue-N and CEConv-N architectures. We trained Hue-3, CEConv-3, Hue-4 and CEConv-4 on CIFAR-10 and used the resulting representations to sort the automobile class by color. Figures 11, 12, 13, and 14 show the automobile class sorted by Hue-3 (ours), CEConv-3, Hue-4 (ours), and CEConv-4 respectively. Figures 11, and 12 are qualitatively similar; however, Figures 13 and 14 are not. In particular, notice that there are blue images in the green band (around the center) of Figure 14, but not in Figure 13.
@@ C.3 Color sorting on CIFAR-10
D Datasets
D.1 Hue-shift MNIST
@@ D.2 Hue-shift 3D Shapes
D.3 small NORB
@@ D.4 Tiny-ImageNet
E Implementation and training details
In this section we provide architectural details for our equivariant networks, and training details for each experiment. Our source code is publicly available online at https://github.com/CAB-Lab-Princeton/Learning-Color-Equivariant-Representations.
@@ E.2 Hue-shift 3D Shapes
Following (Wiles et al., 2022), we train the ResNet architectures for 100k iterations with a batch size of 128. We optimize over a cross-entropy loss using SGD with a learning rate of $10^{-2}$. Images are down-sampled by a factor of 2 to train the CNN architectures. CNN architectures are trained as described in Section E.1.
E.3 Camelyon17
@@ E.11 ImageNet
F Additional Results
F.1 Luminance shift small NORB
We demonstrate improved generalization to luminance-shifts compared to ResNet18 (He et al., 2016) on the small NORB classification dataset (LeCun et al., 2004). The small NORB dataset consist of 5 categories of objects under different lighting conditions. Variation in images reflects variation in the location and strength of the lighting. Additional details are provided in Appendix D.3.
@@ F.1 Luminance shift small NORB
Networks A/A A/B A/C Params
ResNet18 8.32 (0.86) 37.70 (1.51) 33.88 (5.11) 11.2M
Lum-3* 7.04 (0.65) 18.45 (4.83) 25.14 (1.07) 11.1M
F.2 ImageNet
We include classification results on the ImageNet dataset (Deng et al., 2009). Additional details are provided in Appendix E.
@@ F.2 ImageNet
F.3 Foreground-Background Segmentation